# Information-Theoretic Thresholds for the Alignments of Partially Correlated Graphs

### v1

@@ 3.1 Proof of Proposition 1
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)$
$\leq$ $\exp\left(mh\left(\frac{k}{m}\right)-\frac{N_{k}p_{11}\eta^{2}}{2}\right)\leq\exp\left(-mh\left(\frac{k}{m}\right)\right).$
When $k=m$, since $N_{k}=\frac{mk}{2}\left(2-\frac{k+1}{m}\right)\geq\frac{mk}{3}$, we have
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right)\leq\exp\left(-\frac{\log n}{6}\right).$
Pick $c_{1}(\delta)=100\vee\frac{200h(1-\delta)}{1-\delta}$. We then verify the condition in Lemma 3: $\eta\leq\frac{\gamma}{4(1+\gamma)}$. Since $p^{2}\phi(\gamma)\leq p^{2}\gamma(1+\gamma)\leq 1$, we get $m\geq c_{1}(\delta)\log n$. Therefore,
$\eta$ $\leq\sqrt{\frac{8h(1-\delta)}{(1-\delta)mp_{11}}}\cdot\mathbbm{1}_{k\leq m-1}+\sqrt{\frac{\log n}{m^{2}p_{11}}}\cdot\mathbbm{1}_{k=m}\leq\left(\sqrt{\frac{8h(1-\delta)}{1-\delta}}\vee\frac{1}{\sqrt{c_{1}(\delta)}}\right)\frac{1}{\sqrt{mp_{11}}}$
$\leq\left(\sqrt{\frac{8h(1-\delta)}{(1-\delta)c_{1}(\delta)}}\vee\frac{1}{c_{1}(\delta)}\right)\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\leq\frac{1}{5}\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}.$
Recall the assumption stated in Section 1.1, where it’s asserted that $p\geq n^{-1}$, thereby implying $\log(1+\gamma)\leq\log n$. When $\gamma>10$, $\eta\leq\frac{1}{5}\leq\frac{\gamma}{4(1+\gamma)}$. When $\gamma\leq 10$, since $\log(1+x)-\frac{x}{1+x}-x^{2}\leq 0$ for any $x>0$, $\sqrt{\frac{1}{mp_{11}}}\leq\sqrt{\frac{\log(1+\gamma)-\gamma/(1+\gamma)}{\log n}}\leq\frac{\gamma}{\sqrt{\log n}}\leq\frac{\gamma}{4(1+\gamma)}$. Therefore, we obtain $\eta\leq\frac{\gamma}{4(1+\gamma)}$. By Lemma 3, $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$. Applying Lemma 2, we derive (10). Combining this with (19) in Lemma 3, we obtain
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=$ $n^{3k}\exp\left\{-\frac{N_{k}p^{2}}{2}\phi\left[(1+\gamma)(1-\eta)-1\right]\right\}\leq n^{3k}\exp\left(-\frac{N_{k}p^{2}}{8}\phi(\gamma)\right).$
Sum over $k\geq(1-\delta)m$, since $N_{k}\geq\frac{km}{3}$, we obtain
$\sum_{k=\delta m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $\sum_{k=\delta m}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $\exp\left(-\frac{\log n}{6}\right)+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right]+\sum_{k=(1-\delta)m}^{m}\left[n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\right]^{k}$
$\leq$ $\exp\left(-\frac{\log n}{6}\right)+\frac{\exp\left[-(1-\delta)m\log n\right]}{1-\exp(-\log n)}+\sum_{k=(1-\delta)m}^{m-1}\exp\left[-mh\left(\frac{k}{m}\right)\right].$
Combining this with Lemma 4, $\sum_{k=(1-\delta)m}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]\leq(\log n)^{-1+o(1)}$. We finish the proof.
3.2 Proof of Proposition 2
Let $\tau_{k}=N_{k}p_{11}(1-\eta)$ with $\eta=\frac{\gamma}{4(1+\gamma)}$, by (9) and the Chernoff bound (21),
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]\leq\binom{m}{k}\exp\left(-\frac{N_{k}p_{11}\eta^{2}}{2}\right).$
By Lemma 3, $\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}=(1+\gamma)(1-\eta)>1$. Applying Lemma 2, we derive (10). Combining this with (19) in Lemma 3, we obtain
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $n^{3k}\exp\left(-\frac{\tau_{k}}{2}\log\left(\frac{\tau_{k}}{|\mathcal{E}_{\pi}|p^{2}}\right)+\frac{\tau_{k}}{2}-\frac{|\mathcal{E}_{\pi}|p^{2}}{2}\right)$
$=$ $n^{3k}\exp\left\{-\frac{N_{k}p^{2}}{2}\phi\left[(1+\gamma)(1-\eta)-1\right]\right\}\leq n^{3k}\exp\left[-\frac{N_{k}p^{2}}{8}\phi(\gamma)\right].$
Sum over $k\geq 1$, since $N_{k}\geq\frac{km}{3}$ and $\tbinom{m}{k}\leq m^{k}$, we obtain
$\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]$
$\leq$ $\sum_{k=1}^{m}\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}[F_{\pi}])<\tau_{k}\right\}\right]+\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi})-\mathsf{{e}}({\mathcal{H}}_{\pi}[F_{\pi}])\geq\tau_{k}\right\}\right]$
$\leq$ $\sum_{k=1}^{m}\left[m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\right]^{k}+\left[n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\right]^{k}.$
Pick the universal constant $C=400$. Recall that $\phi(\gamma)=(1+\gamma)\log(1+\gamma)-\gamma$. When $\gamma\leq 1$, since $\phi(\gamma)\leq\frac{\gamma^{2}}{1+\gamma}$, we obtain $p^{2}\phi(\gamma)\leq 16p_{11}\eta^{2}$. Therefore, $m\geq\frac{400\log n}{p^{2}\phi(\gamma)}$ implies $m\geq\frac{12\log m}{p_{11}\eta^{2}}$. When $\gamma>1$, since $\gamma\leq 32(1+\gamma)\left[\frac{\gamma}{4(1+\gamma)}\right]^{2}$, we obtain $p^{2}\gamma\leq 32p_{11}\eta^{2}$. Since $m\geq\frac{400\log(1/p^{2}\gamma)}{p^{2}\gamma}$, we have $m\geq\frac{384\log m}{p^{2}\gamma}\geq\frac{12\log m}{p_{11}\eta^{2}}$. Thus we get $m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq\exp(-\log m)$. When $m\geq\frac{400\log n}{p^{2}\phi(\gamma)}$, we get $n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\leq\exp(-\log n)$. Therefore, when $m\geq 400\left(\frac{\log(1/p^{2}\gamma)}{p^{2}\gamma}\vee\frac{\log n}{p^{2}\phi(\gamma)}\right)$, we have
$m\exp\left(-\frac{mp_{11}\eta^{2}}{6}\right)\leq\exp(-\log m),n^{3}\exp\left(-\frac{mp^{2}\phi(\gamma)}{24}\right)\leq\exp(-\log n).$
Therefore, $\sum_{k=1}^{m}\mathbb{P}\left[d(\pi^{*},\hat{\pi})=k\right]\leq\frac{\exp(-\log m)}{1-\exp(-\log m)}+\frac{\exp(-\log n)}{1-\exp(-\log n)}$. We finish the proof.
@@ 4 Impossibility results
  • Choose the uniform prior on $\pi^{*}$ over ${\mathcal{M}}$ and upper bound the mutual information $I(\pi^{*};G_{1},G_{2})$. Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definition 3. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$, then
@@
    The impossibility results follows if $I(\pi^{*};G_{1},G_{2})\leq c\log|{\mathcal{M}}|$ for some small constant $c$.
Let ${\mathcal{M}}_{\delta}$ denote a packing set under the overlap threshold $\delta$. The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW22, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then we have
@@
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size $m-k$ and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\tbinom{m}{k}\tbinom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
@@
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bound the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via generalized Fano’s inequality [BHM12, Lemma 20]. The following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2.
@@
For any $\delta\in(0,1)$, if $m\leq\frac{c\log n}{p^{2}\phi(\gamma)}$, then for any estimator $\hat{\pi}$,
@@
For any $\pi$ with domain $S$ and range $T$ such that $|S|=|T|=m$, arbitrarily pick a bijection $\sigma:V(G_{1})\mapsto V(G_{2})$ such that $\sigma|_{S}=\pi$. Then, the conditional distribution $\mathcal{P}_{G_{1},G_{2}|\pi}$ can be factorized into
$\mathcal{P}_{G_{1},G_{2}|\pi}=\prod_{e\in\tbinom{S}{2}}P(e,\pi(e))\prod_{e\in\tbinom{V(G_{1})}{2}\backslash\tbinom{S}{2}}Q(e,\sigma(e)),$
where $P\sim\mathrm{Bern}(p,p,\rho)$ and $Q\sim\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ in (11) to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\tbinom{S}{2}}Q(e,\pi(e))\prod_{e\in\tbinom{V(G_{1})}{2}\backslash\tbinom{S}{2}}Q(e,\sigma(e)).$
@@
$D\left(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}\right)=\tbinom{m}{2}D(P\|Q)$
@@
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\tbinom{m}{2}p^{2}\phi(\gamma).$ (13)
@@
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\tbinom{m}{2}p^{2}\phi(\gamma)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
@@
Proposition 4 (Lower bound for exact recovery).
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists constant $c_{3}$ only depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)$,
@@
When $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\asymp n$, by (14), exact recovery is impossible, even when $m=n$. Next we consider the regime that $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\lesssim n$. When $\gamma\leq 1$, we have $p^{2}(\gamma\wedge\gamma^{2})=p^{2}\gamma^{2}\asymp p^{2}\phi(\gamma)$, and thus

### v2

@@ 3.2 Gaussian Wigner model
$\mathbb{P}\left[\bigcup_{\pi\in{\mathcal{T}}_{k}}\left\{\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f})-\mathsf{{e}}({\mathcal{H}}_{\pi^{*}}^{f}[F_{\pi}])<\tau_{k}\right\}\right]$ $\leq\binom{m}{k}\mathbb{P}\left[\sum_{i=1}^{N_{k}}-\frac{1}{2}(A_{i}-B_{i})^{2}<\tau_{k}\right],$ (17)
the tail follows from the concentration inequality for chi-squared distribution in Lemma 10. For the bad event of noise, applying (8) with the formula of lower-order cumulants (45) and (46) yields the following Lemma, whose proof is deferred to Appendix B.4.
Lemma 4.
In the Gaussian Wigner model, for $f(x,y)=-\frac{1}{2}(x-y)^{2}$, any threshold $\tau_{k}$, and $\pi\in\mathcal{S}_{n,m}$ with $k=d(\pi,\pi^{*})$, we have
$\mathbb{P}\left[\beta_{\mathcal{E}_{\pi}}(\mathcal{H}_{\pi}^{f})\geq\tau_{k}\right]\leq\exp\left(-\frac{\rho\tau_{k}}{4(1-\rho)}-\frac{|\mathcal{E}_{\pi}|}{4}\log\left(\frac{1}{1-\rho}\right)+\frac{k}{8}\log\left(\frac{1}{1-\rho}\right)\right).$
The following proposition provides a sufficient condition on $m$ for exact recovery in the Gaussian Wigner model when $11-e^{-12}<\rho<$ 1, whose proof is deferred to Appendix A.4.
Proposition 4.
When $11-e^{-12}<\rho<$ 1, there exists a universal constant $C_{4}>0$ such that, when $m\geq C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right)$, for any $\pi^{*}\in\mathcal{S}_{n,m}$, the estimator in (6) with $f(x,y)=-\frac{1}{2}(x-y)^{2}$ satisfies
$\mathbb{P}\left[\hat{\pi}\neq\pi^{*}\right]\leq\frac{2}{n-1}.$
In view of Propositions 3 and 4, we note that $\rho^{2}\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $0<\rho\leq 1-e^{-12}$, and $\log\left(\frac{1}{1-\rho}\right)\asymp\log\left(\frac{1}{1-\rho^{2}}\right)$ when $11-e^{-12}<\rho<$ 1. Then, there exists a universal constant $C_{2}$ such that,
$C_{2}\left(\frac{\log n}{\log(1/(1-\rho^{2}))}\vee 1\right)\geq\begin{cases}\frac{C_{3}\log n}{\rho^{2}},&0<\rho\leq 1-e^{-12},\\ C_{4}\left(\frac{\log n}{\log(1/(1-\rho))}\vee 1\right),&1-e^{-12}<\rho<1.\end{cases}$
Therefore, we prove the possibility results in Theorem 3.
Remark 3.
While the proofs for the possibility results in the Erdős-Rényi and Gaussian Wigner models share a similar structure, they also exhibit several key differences. Following the intuition of MLE in (9), we analyze two different estimators for the Gaussian Wigner model. Furthermore, by applying different concentration inequalities, we observe a significant difference in the choice of $\tau_{k}$. Specifically, we set $\tau_{k}=\Theta\left(\mathbb{E}\left[\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}\right)-\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}[F_{\pi}]\right)\right]\right)$ for both models with $\rho=1-\Omega(1)$. In contrast, we choose $\tau_{k}=\omega\left(\mathbb{E}\left[\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}\right)-\mathsf{{e}}\left(\mathcal{H}_{\pi^{*}}^{f}[F_{\pi}]\right)\right]\right)$ for the Gaussian model with $\rho=1-o(1)$. The thresholds are chosen to balance the error probabilities arising from the bad events of signal and noise, ultimately achieving the optimal rates for the correlated size $m$.
Remark 4.
Our estimator (6) requires knowledge of the cardinality $|S^{*}|=m$. When $m$ is unknown, the objective function in (6) will increase monotonically with $m$. To address this, we can introduce a penalty term for $m$ in the estimator to identify the correct size, following an approach similar to the MLE in the new model. This will be discussed in more detail in a forthcoming working paper.
@@ 4 Impossibility results
  • Choose the uniform prior on $\pi^{*}$ over ${\mathcal{M}}$ and upper bound the mutual information $I(\pi^{*};G_{1},G_{2})$. Given $\pi^{*}$, the conditional distribution of the observed graphs $(G_{1},G_{2})$ is specified in Definition 4. For the mutual information, let $\mathcal{P}$ denote the joint distribution of $(G_{1},G_{2})$ and $\mathcal{Q}$ be any distribution over $(G_{1},G_{2})$. Then,
@@
    The impossibility results follow if $I(\pi^{*};G_{1},G_{2})\leq c\log|{\mathcal{M}}|$ for some small constant $c$.
Let ${\mathcal{M}}_{\delta}$ denote a packing set under the overlap threshold $\delta$. The size of ${\mathcal{M}}_{\delta}$ follows from the standard volume argument [PW25, Theorem 27.3]. For $r\in[m]$, let $B(\pi,r)\triangleq\{\pi^{\prime}:d(\pi,\pi^{\prime})\leq r\}$ denote the ball of radius $r$ centered at $\pi$. Then, we obtain
@@
It remains to evaluate the cardinality of $\mathcal{S}_{n,m}$ and upper bound the volume of the ball under our distance metric $d$. It is straightforward to obtain that $|\mathcal{S}_{n,m}|=\binom{n}{m}^{2}m!.$ Let $k=\delta m$. Note that all elements from $B(\pi,m-k)$ have at least $k$ common mappings. To upper bound $|B(\pi,m-k)|$, we first choose $k$ elements from the domain of $\pi$ and map to the same value as $\pi$, and the remaining domain and range of size $m-k$ and the mapping are selected arbitrarily. We get $|B(\pi,m-k)|\leq\binom{m}{k}\binom{n-k}{m-k}^{2}(m-k)!.$ Consequently,
@@
where we use the inequalities that $(\frac{n}{k})^{k}\leq\binom{n}{k}<(\frac{en}{k})^{k}$ and $k!\geq(k/e)^{k}$. Fano’s method provides a lower bound on the Bayesian risk when $\pi$ is uniformly distributed over ${\mathcal{M}}_{\delta}$, which further lower bounds the minimax risk. The above argument also yields a lower bound when $\pi$ is uniform over $\mathcal{S}_{n,m}$ via generalized Fano’s inequality [BHM12, Lemma 20]. For the Erdős-Rényi model, the following propositions provide lower bounds for $m$ for partial recovery and exact recovery, and thus prove the lower bounds in Theorems 1 and 2. For the Gaussian Wigner model, the Proposition 5 provides a lower bound for $m$ for partial recovery, which implies that $m\leq\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}$ is sufficient for the lower bound since $m<1$ is impossible for partial recovery. Consequently, we finish the proof of impossibility results in Theorem 3 since $\frac{c\log n}{\log\left(1/(1-\rho^{2})\right)}\vee\frac{1}{2}\asymp\frac{\log n}{\log\left(1/(1-\rho^{2})\right)}\vee 1$.
@@
In the Erdős-Rényi model, for any $\delta\in(0,1)$, if $m\leq\frac{c\log n}{p^{2}\phi(\gamma)}$, then for any estimator $\hat{\pi}$,
@@
In the Gaussian Wigner model, for any $\delta\ in(0,1)$, if $m\leq\frac{c\log n}{\log(1/(1-\rho^{2}))}$, then for any estimator $\hat{\pi}$,
$\mathbb{P}\left[\mathrm{overlap}(\hat{\pi},\pi^{*})<\delta\right]\geq 1-\frac{c}{2\delta}.$
@@
In the Erdős-Rényi model, for any $\pi$ with domain $S$ and range $T$ such that $|S|=|T|=m$, we arbitrarily pick a bijection $\sigma:V(G_{1})\mapsto V(G_{2})$ such that $\sigma|_{S}=\pi$. Then, the conditional distribution $\mathcal{P}_{G_{1},G_{2}|\pi}$ can be factorized into
$\mathcal{P}_{G_{1},G_{2}|\pi}=\prod_{e\in\binom{S}{2}}P(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)),$
where $P=\mathrm{Bern}(p,p,\rho)$ and $Q=\mathrm{Bern}(p,p,0)$. Pick $\mathcal{Q}$ in (18) to be an auxiliary null model under which $G_{1}$ and $G_{2}$ are independent with the same marginal as $\mathcal{P}$. Then, $\mathcal{Q}_{G_{1},G_{2}}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\binom{S}{2}}Q(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)).$
@@
$D\left(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}}\right)=\binom{m}{2}D(P\|Q)$
@@
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq 25\binom{m}{2}p^{2}\phi(\gamma).$ (20)
@@
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{25\binom{m}{2}p^{2}\phi(\gamma)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{13c}{\delta},$
where $\pi^{*}$ is uniformly distributed over ${\mathcal{S}}_{n,m}$.
In the Gaussian Wigner model, let $P=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&\rho\\ \rho&1\end{pmatrix}\right)$ and $Q=\mathcal{N}\left(\begin{pmatrix}0\\ 0\end{pmatrix},\begin{pmatrix}1&0\\ 0&1\end{pmatrix}\right)$. We can similarly pick the auxiliary null model $\mathcal{Q}$ under which $G_{1}$ and $G_{2}$ are independent. Then, $\mathcal{Q}$ can be factorized into
$\mathcal{Q}_{G_{1},G_{2}}=\prod_{e\in\binom{S}{2}}Q(e,\pi(e))\prod_{e\in\binom{V(G_{1})}{2}\backslash\binom{S}{2}}Q(e,\sigma(e)).$
Applying Lemma 7, we obtain
$\max_{\pi}D(\mathcal{P}_{G_{1},G_{2}|\pi}\|\mathcal{Q}_{G_{1},G_{2}})\leq\binom{m}{2}D(P\|Q)\leq\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right).$ (21)
$\mathbb{P}\left[\mathrm{overlap}(\pi^{*},\hat{\pi})<\delta\right]\geq 1-\frac{\frac{1}{2}\binom{m}{2}\log\left(\frac{1}{1-\rho^{2}}\right)}{\delta m\log\left(\frac{\delta n}{e^{3}}\right)}\geq 1-\frac{c}{2\delta},$
@@
Proposition 6 (Erdős-Rényi model, lower bound for exact recovery).
For any $c\in(0,1)$ and any estimator $\hat{\pi}$, there exists a constant $c_{3}$ depending on $c$ such that, when $m\leq c_{3}\left(\frac{\log n}{p^{2}\phi(\gamma)}\vee\frac{1}{p^{2}\gamma}\log\left(\frac{1}{p^{2}\gamma}\right)\right)$,
@@
When $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\asymp n$, by (22), exact recovery is impossible, even when $m=n$. Next, we consider the regime that $\frac{1}{p^{2}(\gamma\wedge\gamma^{2})}\lesssim n$. When $\gamma\leq 1$, we have $p^{2}(\gamma\wedge\gamma^{2})=p^{2}\gamma^{2}\asymp p^{2}\phi(\gamma)$, and thus